OpenStudy (anonymous):

How many license plates can be made which have the form ##LLL? 1,757,600 175,600 3,840,500 2,947,000

OpenStudy (anonymous):

Using fundamental principle of counting: a. if repetitions are allowed: N = 10 * 10 * 26 * 26 * 26 (10 digits, 26 letters) N = 1,757,600 b. if repetitions are not allowed: N = 10 * 9 * 26 * 25 * 24 (deduct 1 for every used character) N = 1,404,000 i have answered in yahoo.. so never mind thats the same...